OpenStudy (kenneady):

(26 div 13) x (-7)+14-4 pls anyone

OpenStudy (wckeller):

Let's start by working in the parentheses... What is 26/13=?

OpenStudy (kenneady):

2???

OpenStudy (aveline):

We have: \[\frac{ 26 }{ 13 } \times (-7) + 14 - 4\] You have already simplified 26/13 \[2 \times (-7) +14 -4\]

OpenStudy (aveline):

Use PEMDAS. First, multiply 2 and -7. Then add 14. Finally, subtract 4.

OpenStudy (kenneady):

is 24 right???

OpenStudy (aveline):

2 x -7 = -14 Don't forget the negative!

OpenStudy (kenneady):

soo negative 24?? sorry im really bad at math

OpenStudy (aveline):

2 x -7 = -14 Now we add 14 -14 + 14 = 0 Finally, we subtract 4 0-4 = -4 Our answer is negative 4, or -4

OpenStudy (wckeller):

Sorry @kenneady I got wrapped up in other problems while I was waiting for your response and forgot to come back!

OpenStudy (kenneady):

it fine !!!

OpenStudy (kenneady):

and thank you so uch!!!!!!!!!